In ∆ABC, C=30 degrees, c=22, and b=42. Find B.

Answer by Nate(3500)

Use Law of Sines:
Sin(B)/b = Sin(C)/c
Sin(B)/42 = Sin(30pi/180)/22
Sin(B) = 42*Sin(pi/6)/22
Sin(B) = 21*Sin(pi/6)/11
about 73 degrees